# Cuntz–Pimsner algebras of partial automorphisms twisted by vector bundles I: Fixed point algebra, simplicity and the tracial state space

### v2

@@ 3.5. Fibered structure of the fixed point algebra
We write $\mathcal{E}$ for the Hilbert $C_{0}(X)$-module $\Gamma_{0}(\mathcal{V},\theta)$, where $\theta:U\to V$ is a partial automorphism and $\mathcal{V}$ is a vector bundle over $U$. Section 3.1 shows that both $\mathcal{E}$ and $\mathcal{K}(\mathcal{E})$ naturally fiber over $U$, and hence over $X$: More precisely, the fiber of $\mathcal{E}$ at a point $x\in U$ is isomorphic to $\mathbb{C}^{d(x)}$, where $d(x)$ is the rank of the vector bundle at the point $x$. For $x\in X\backslash U$ the fiber is $\{0\}$, the trivial Hilbert space. The fiber of $\mathcal{K}(\mathcal{E})$ at a point $x\in U$ is $\mathcal{K}(\mathcal{E}_{x})\cong M_{n_{x}}(\mathbb{C})$, and at a point $x\in X\backslash U$ it is the trivial $C^{*}$-algebra, which we also denote by $\{0\}$.
@@
from Proposition 3.20, where we have given names to the maps now. Lemmas 3.22 and 3.23 together with [29, Lemma 5.10] imply that $\iota^{n}(B_{n}^{\prime})B_{n+1}$ equals $B_{n+1}$. Theorem 2.4 in [35] then shows that the diagram is a pushout. From Lemma 3.24 we obtain that the diagram
@@
If the rank of the vector bundle is constant and equal to $d\in\mathbb{N}$, then the fiber over a point $x\in\bigcap_{k\geq 0}D_{k}$ is given by the UHF algebra $\mathcal{U}_{d^{\infty}}$.
@@
Assume that $X$ is compact and has finite covering dimension, $\theta:X\to X$ has domains $U=V=X$ and that the vector bundle has constant rank $d$. Then the fixed point algebra is trivial as a $C(X)$-algebra, meaning that $\mathcal{O}(\mathcal{E})^{0}$ is isomorphic to $C(X)\otimes\mathcal{U}_{d^{\infty}}$ as $C(X)$-algebras.
@@ 4.2. Restricting the $C^{*}$-correspondence
This section compiles a number of results which are related to restricting the partial automorphism and hence the $C^{*}$-correspondence in different ways, and which we will need in later chapters.
@@ 4.3. Ideal structure
Take $\varepsilon>0$ and $a\in\mathcal{O}(\mathcal{E})$. According to Proposition 4.9, there exists an open set $W$ whose closure is compact and contained in $U$, and such that $\left\|a-b\right\|<\varepsilon$ for some $b$ in $\mathcal{O}(\mathcal{E}^{(W)})$. The algebraic direct sum over all spectral subspaces $\mathcal{O}(\mathcal{E}^{(W)})^{n}$ is dense in $\mathcal{O}(\mathcal{E}^{(W)})$. Therefore, by Proposition 3.18, we can assume that $b$ has the form $\sum_{n=-N}^{N}g_{n}b_{n}$ for $g_{n}\in C_{0}(D^{(W)}_{n})$, $b_{n}\in\mathcal{O}(\mathcal{E}^{(W)})^{n}$, and some $N\in\mathbb{N}$. We can further assume that the fibers of $g_{0}b_{0}$ in the fixed point algebra (see Theorem 3.27) vanish outside of some compact set $C$ contained in $U$.
@@ A Radon probability measure $\mu$ on $X$ is called $\theta$-invariant if
Up to a metrizability assumption which can easily be removed, this is a special case of [5, Theorem 3.15]. The existence also follows from [41, Proposition 2.7] together with the fact that $\mathbb{Z}$ as an abelian group is supramenable. ∎
@@
If $\mathcal{V}=[E,p,U]$ is not a line bundle, namely if $\operatorname{rank}p^{-1}(x)>1$ for some $x\in U$, then $\mathcal{O}(\mathcal{E})$ has no faithful tracial state. If $X$ is compact and $\mathcal{V}$ is a line bundle then $T(\mathcal{O}(\mathcal{E}))\neq 0$.
@@
Let $x_{0}$ be a point in $U$ such that $\operatorname{rank}p^{-1}(x_{0})=M$ with $M>1$. Let $W$ be an open set containing $x_{0}$ such that $h:W\times\mathbb{C}^{M}\to\mathcal{V}|_{W}$ is a local trivialization of $\mathcal{V}$. We can assume that $W$ is contained in $U$.
Define $\xi_{k}(x):=h(x,\gamma^{\frac{1}{2}}(x)e_{k})$, where $e_{1},...,e_{M}$ is the standard orthonormal basis of $\mathbb{C}^{M}$. The $\xi_{k}$ lie in $\mathcal{V}$ because $\gamma$ is supported in $W$ which is contained in $U$. Then
$\xi_{k}^{*}\xi_{k}=\langle\xi_{k},\xi_{k}\rangle_{\mathcal{E}}=\gamma\quad\text{for }1\leq k\leq M.$
On the other hand, for every $\eta$ in $\mathcal{E}$ a short calculation shows $\sum_{k=1}^{M}\xi_{k}\langle\xi_{k},\eta\rangle_{\mathcal{E}}=\gamma\eta.$ Hence $\sum_{k=1}^{M}\xi_{k}\xi_{k}^{*}=\gamma$ which implies
$\tau(\gamma)$ $=\tau\left(\sum_{k=1}^{M}\xi_{k}\xi_{k}^{*}\right)=\sum_{k=1}^{M}\tau(\xi_{k}\xi_{k}^{*})$
$=\sum_{k=1}^{M}\tau(\xi_{k}^{*}\xi_{k})=\sum_{k=1}^{M}\tau(\gamma)=M\tau(\gamma).$
Since $M>1$ this implies $\tau(\gamma)=0$. We obtain that $\tau$ is not faithful.
Now assume that $\operatorname{rank}p^{-1}(x)=1$ for all $x\in U$. Let $\mu$ be an $\theta$-invariant probability measure on $X$ in the sense of Definition 4.19. Let $\tau_{\mu}$ be the associated state as in Remark 4.20. We claim that $\tau_{\mu}\circ\Phi$ is tracial.
@@
For $X$ compact and $\mathcal{V}=[E,p,U]$ a line bundle, the map from $M^{1}_{\theta}(X)$ to $T(\mathcal{O}(\mathcal{E}))$ sending $\mu$ to $\tau_{\mu}\circ\Phi$ is an affine homeomorphism.
@@
Let $(\xi_{i})_{i=1}^{m}$ be the frame for $\Gamma(\mathcal{V})$ from Lemma 2.6. Note that the variable $k$ is no longer required since the vector bundle is a line bundle. We have
$\sum_{i=1}^{N}\xi_{i}\xi_{i}^{*}=\sum_{i=1}^{N}\xi_{i}^{*}\xi_{i}=1.$
For any $f\in C_{0}(V)$ we obtain
$\tau(f)$ $=\tau\left(f\sum_{i=1}^{N}\xi_{i}\xi_{i}^{*}\right)=\sum_{i=1}^{N}\tau(f\xi_{i}\xi_{i}^{*})$
$=\sum_{i=1}^{N}\tau(\xi_{i}(f\circ\theta)\xi_{i}^{*})=\sum_{i=1}^{N}\tau(\xi_{i}^{*}\xi_{i}(f\circ\theta))$
@@
which shows that $\tau|_{C(X)}$ is $\theta$-invariant.

### v3

@@ 3.5. Fibered structure of the fixed point algebra
We write $\mathcal{E}$ for the Hilbert $C_{0}(X)$-module $\Gamma_{0}(\mathcal{V},\theta)$, where $\theta:U\to V$ is a partial automorphism and $\mathcal{V}$ is a vector bundle over $U$. Section 3.1 shows that both $\mathcal{E}$ and $\mathcal{K}(\mathcal{E})$ naturally fiber over $U$, and hence over $X$: More precisely, the fiber of $\mathcal{E}$ at a point $x\in U$ is isomorphic to $\mathbb{C}^{d(x)}$, where $d(x)$ is the rank of the vector bundle at the point $x$. For $x\in X\backslash U$ the fiber is $\{0\}$, the trivial Hilbert space. The fiber of $\mathcal{K}(\mathcal{E})$ at a point $x\in U$ is $\mathcal{K}(\mathcal{E}_{x})\cong M_{d(x)}(\mathbb{C})$, and at a point $x\in X\backslash U$ it is the trivial $C^{*}$-algebra, which we also denote by $\{0\}$.
@@
from Proposition 3.21, where we have given names to the maps now. Lemmas 3.23 and 3.24 together with [29, Lemma 5.10] imply that $\iota^{n}(B_{n}^{\prime})B_{n+1}$ equals $B_{n+1}$. Theorem 2.4 in [35] then shows that the diagram is a pushout. From Lemma 3.25 we obtain that the diagram
@@
If the rank of the vector bundle is constant and equal to $d\in\mathbb{N}$, then the fiber over a point $x\in\bigcap_{k\geq 0}D_{k}$ is given by the UHF algebra $M_{d^{\infty}}$.
@@
Assume that $X$ is compact and has finite covering dimension, $\theta:X\to X$ has domains $U=V=X$ and that the vector bundle has constant rank $d$. Then the fixed point algebra is trivial as a $C(X)$-algebra, meaning that $\mathcal{O}(\mathcal{E})^{0}$ is isomorphic to $C(X)\otimes M_{d^{\infty}}$ as $C(X)$-algebras.
@@ 4.2. Restricting the $C^{*}$-correspondence
This section compiles a number of results which are related to restricting the partial automorphism and hence the $C^{*}$-correspondence in different ways, and which we will need in later chapters. We write $\mathcal{E}$ for the $C^{*}$-correspondence $\Gamma_{0}(\mathcal{V},\theta)$ over $C_{0}(X)$, where $\theta:U\to V$ is a partial automorphism on a locally compact second countable Hausdorff space $X$, and $\mathcal{V}$ is a vector bundle over $U$.
@@ 4.3. Ideal structure
Take $\varepsilon>0$ and $a\in\mathcal{O}(\mathcal{E})$. According to Proposition 4.9, there exists an open set $W$ whose closure is compact and contained in $U$, and such that $\left\|a-b\right\|<\varepsilon/2$ for some $b$ in $\mathcal{O}(\mathcal{E}^{(W)})$. The algebraic direct sum over all spectral subspaces $\mathcal{O}(\mathcal{E}^{(W)})^{n}$ is dense in $\mathcal{O}(\mathcal{E}^{(W)})$. Therefore, by Proposition 3.18, we can assume that $b$ has the form $\sum_{n=-N}^{N}g_{n}b_{n}$ for $g_{n}\in C_{0}(D^{(W)}_{n})$, $b_{n}\in\mathcal{O}(\mathcal{E}^{(W)})^{n}$, and some $N\in\mathbb{N}$. We can further assume that the fibers of $g_{0}b_{0}$ in the fixed point algebra (see Theorem 3.28) vanish outside of some compact set $C$ contained in $U$.
@@ A Radon probability measure $\mu$ on $X$ is called $\theta$-invariant if
Up to a metrizability assumption which can be removed, this is a special case of [5, Theorem 3.15]. The existence also follows from [41, Proposition 2.7] together with the fact that $\mathbb{Z}$ as an abelian group is supramenable. ∎
@@
If $X$ is compact and $\mathcal{V}$ is a line bundle then $T(\mathcal{O}(\mathcal{E}))\neq 0$.
@@
Let $\mu$ be an $\theta$-invariant probability measure on $X$ in the sense of Definition 4.19. Let $\tau_{\mu}$ be the associated state as in Remark 4.20. We claim that $\tau_{\mu}\circ\Phi$ is tracial.
@@
If $X$ is compact, $\theta$ is free, and $\mathcal{V}$ is a line bundle, then the map from $M^{1}_{\theta}(X)$ to $T(\mathcal{O}(\mathcal{E}))$ sending $\mu$ to $\tau_{\mu}\circ\Phi$ is an affine homeomorphism.
@@
Using an atlas for $\mathcal{V}$ and a suitable partition of unity, we can construct a countable frame $(\xi_{i})_{i=1}^{\infty}$ of $\Gamma_{0}(\mathcal{V},\theta)$ similarly to Lemma 2.6. It has the property
$\sum_{i=1}^{\infty}\xi_{i}^{*}\xi_{i}=1.$
Under the left action any $f\in C_{0}(V)$ acts on $\Gamma_{0}(\mathcal{V},\theta)$ as a compact operator, which implies
$f\sum_{i=1}^{\infty}\xi_{i}\xi_{i}^{*}=f.$
Thus
$\tau(f)$ $=\tau\left(f\sum_{i=1}^{\infty}\xi_{i}\xi_{i}^{*}\right)=\sum_{i=1}^{\infty}\tau(f\xi_{i}\xi_{i}^{*})$
$=\sum_{i=1}^{\infty}\tau(\xi_{i}(f\circ\theta)\xi_{i}^{*})=\sum_{i=1}^{\infty}\tau(\xi_{i}^{*}\xi_{i}(f\circ\theta))$
@@
where in the second last step we have used that $\tau$ is a trace. This shows that $\tau|_{C(X)}$ is $\theta$-invariant.